OpenStudy (anonymous):

An algebra test contains 38 problems. some of the problems are worht 2 points each and the rest are worth 3 poins each. A perfect score is 100 points. How many problems are worth 2 points?

OpenStudy (aaronq):

you need 2 equations: 38 problems, 2 types (1) 38=x+y these have to sum up to 100, so (2) 3x + 2y = 100 Now substitute one into the other (I suggest (1) into (2)) and solve.

OpenStudy (anonymous):

so whats the answer? i wana make sure i got the right one

OpenStudy (aaronq):

post what you did if you want me to check whether it's right or not

OpenStudy (anonymous):

3x+2(1)=100 3x+2=100 -2 -2 3x=98 98 divided 3 is 32.6

OpenStudy (aaronq):

why did you plug in 1 into y?

OpenStudy (anonymous):

didnt u say that?

OpenStudy (aaronq):

i meant equation (1), not the number one...

OpenStudy (anonymous):

ohhh okay let me trye again

OpenStudy (aaronq):

okay

OpenStudy (anonymous):

idk what to do! ikno im suppose to plug it in but how? is this correct: 3x+2(38)=100

OpenStudy (aaronq):

nope, you need to isolate a variable, say x: 38=x+y -> x=38-y then plug that into the second equation

OpenStudy (anonymous):

ohhh so 3(38-y)+2y=100

OpenStudy (aaronq):

yeah thats it, then solve for y. once you know y, you can find x (by using the first equation).

OpenStudy (aaronq):

remember what those variables represent though

OpenStudy (anonymous):

okay thank you soo much do u think u can help me with something else?

OpenStudy (anonymous):

and for the answer i got 14

OpenStudy (aaronq):

thats right

OpenStudy (aaronq):

sure, shoot

OpenStudy (anonymous):

okay so write the function h(x)=-5(x-1)^2+3

OpenStudy (anonymous):

in standard form

OpenStudy (aaronq):

expand this bracket (x-1)^2, then distribute 5 to all the terms within it then simplify

OpenStudy (aaronq):

you should be asking these questions in the math section :P

OpenStudy (anonymous):

im new to this so idknow how to work this website out

OpenStudy (aaronq):

no worries. it's up at the top, click on "Find More Subjects" and a menu will pop up.

OpenStudy (anonymous):

its not letting me it freezes :(

OpenStudy (anonymous):

could u help me with this??

OpenStudy (aaronq):

I can't help you with all of that at the moment, but i can give you a link to get to the math section http://openstudy.com/study#/groups/Mathematics